# and quantum information science

### v1

@@ 5.4 Relation between Nielsen complexity and entanglement: binding complexity
The penalty factors in the definition of Nielsen’s complexity can be attributed to the relative difficulty of applying certain generators. A particularly interesting choice, where analytic results can be obtained, is the case in which our system is split into subsystems, and the penalty factors associated with generators acting within a given subsystem are much smaller than those of generators acting between multiple subsystems. This setting serves as a model to describe distributed quantum computing, where a large computer is divided into small quantum computers (called nodes). The operations between nodes, which are possibly located far away from each other, require significantly more resources than operations within each subsystem.
@@
The limiting case, where only operations acting between subsystems have non-vanishing cost, was referred to as binding complexity in [95]. There, the authors suggested that this notion quantifies the difficulty of distributing entanglement among multiple parties. The authors of Ref. [95] further suggested that the notion of binding complexity could serve as a measure for the robustness of entanglement. Here, one typically has in mind the $n$-party (with $n\geq 2$) generalizations of the W and GHZ states. The GHZ state is separable upon tracing out any subset of the parties, but the W state is not. In this sense, W states possess more robust entanglement. It was argued in [95] that the binding complexity should scale as $O(n)$ for GHZ states, and as $O(n^{2})$ for $W$ states. The paper also presented results for the binding complexity of Gaussian states of free QFT. As we review next, one can get exact results for the binding complexity of spin chains in Nielsen’s framework when the cost of generators acting within each subsystem is small enough [99]. As a bi-product, it was discovered that cost functions satisfying this property relate Nielsen’s complexity to the Rényi min-entropy in a precise way. Moreover, the above cost functions are bounded by the maximal rate of change of a state’s entanglement entropy.
@@
To illustrate these results, let us begin with a formal definition of the binding complexity. Consider splitting a quantum system (for instance, a spin chain) into $m$ subsystems $A_{k=1,2,\ldots,m}\,.$ The binding complexity is the minimal number of gates, each acting on a limited number of subsystems, needed to implement a target unitary [95]. Below, we focus on 2-local gates (which act on, at most, two subsystems each). Within Nielsen’s framework, binding complexity has a close analogue, obtained by choosing a particular cost function on the unitary space. To this aim, we consider a bipartite system whose Hamiltonian (17) decomposes as [^14] We denote the generators of the special unitary group in subsystem $A_{1}$ with index $a\in\{1,\dots,N_{A_{1}}^{2}-1\}$ from the beginning of the alphabet, and the generators in $A_{2}$ with index $i\in\{1,\dots,N_{A_{2}}^{2}-1\}$ from the middle part of the alphabet. We also denote with a sub(super)script $A_{1},A_{2}$ the subsystem where the local generators act.
@@
These penalties impose greater costs on entangling (or non-local) gates than on single-subsystem (or local) operations. Let us substitute the penalties (27) into the cost function (25) and perform the minimization (20) to obtain a complexity norm on the space of quantum states. Two simplifications follow: first, the space of states contains null directions, defined as the loci of states interconnected via local unitaries along which the cost function vanishes. Second, one can integrate out nondynamical degrees of freedom in the geodesic minimization. One obtains a norm $\mathcal{BF}[\lambda_{k},\dot{\lambda}_{k}]$ on the space of states. $\mathcal{BF}$ depends only on a state’s Schmidt coefficients $\lambda_{k}$ and their derivatives $\dot{\lambda}_{k}$.
We can analytically compute the binding state complexity under certain choices of the penalty factors $q_{ai}$ in Eq. (27), and of the parameter $p$ in Eq. (25). A Riemannian cost function provides our first example. Let us write out the binding state complexity associated with the cost function (25), $p=2$, and the penalty factors (27) with $q_{ai}=1$ [99]:
@@
The relationship between binding complexity and Schmidt coefficients leads to not only exact results, but also bounds imposed by other quantum-information quantities. For instance, one can lower-bound the binding state complexity, using the maximal entanglement rate [103, 104, 105]. Denote by $S_{A_{1}}$ the entanglement entropy of subregion $A_{1}$, by $d$ the dimensionality of the smaller subsystem’s Hilbert space, and by $c$ a constant. The $p{=}1$ binding state complexity upper bounds $S_{A_{1}}$ as follows:
@@
Binding complexity also provides lower bounds for other notions of complexity, such as the geometrically local complexity [106]. Further details on this quantity will be discussed around Eq. (142).
Finally, we anticipate a connection between binding complexity and a holographic dual in a black-hole background. The binding complexity has been conjectured to quantify multiboundary wormholes’ interior volumes [95].
@@ 5.5 Complexity in quantum field theories
As we will review later, Nielsen’s complexity experienced a revival following conjectured relationship with gravitational quantities. This connection is rooted in the holographic correspondence, which will be reviewed in Sec. 7. The holographic correspondence deals with QFTs – many-body quantum systems which do not conserve the particle number. To allow for a quantitative comparison with holographic conjectures, it was therefore essential to develop a concrete method for calculating complexity in QFTs. Naturally, the first examples studied were those of free QFTs, first bosonic [84, 85] and then fermionic [88, 89], which we review in subsection 5.5.1. While these calculations are too simple for an exact comparison with chaotic theories in general – and with holography in particular – this approach played a fundamental role in advancing the complexity community and bridging the gap between heuristic arguments and more rigorous calculations. This framework later inspired studies of complexity in more complicated field theories, including weakly interacting theories [91, 107], and conformal field theories, as described in subsections 5.5.2 and 5.5.3 below.
@@ 5.5.1 Free quantum field theories
The calculation of complexity in free field theories is based on Nielsen’s approach (see Sec. 5.1). In this framework, the evolution of quantum states along the circuit is treated continuously, with a cost assigned to each possible evolution. To simplify the problem, one often restricts the circuit evolution to unitary representations of finite-dimensional spaces. For free QFTs, this is typically achieved by restricting the circuits to transitions between Gaussian states. The groups underlying these circuits are the symplectic group for bosonic Gaussian states and the orthogonal group for fermionic Gaussian states.
Like many other quantities in QFT, complexity is expected to diverge. Intuitively, these divergences arises because quantum correlations must be built at arbitrarily short distance scales. Therefore, a regulator must be introduced to handle short distances, or equivalently, large momenta. The exact result depends on the choice of reference state, and in QFT, it is not immediately obvious which reference state to choose. Inspired by quantum computation, the reference state is often chosen to be completely unentangled, meaning it lies outside the Hilbert space of the non-regulated theory, leading to the observed divergences.
@@
where $k_{(ab)}(t)$ is a symmetric matrix, move us between Gaussian states. A natural description of Gaussian states is in terms of their covriance matrices $G^{(ab)}$ and displacement vector $\omega^{a}$ defined as the first and second moments of the phase space variables
@@
where $\hat{\rho}$ is the Gaussian density matrix, and $\Omega^{[ab]}$ encodes the anti-symmetric commutation relations of the phase space variables. The covariance matrices were not the original formulation used to study complexity in bosonic QFT. There, the authors only focused on squeezing operations, and therefore expressing everything in terms of the wavefunction was sufficient. Nevertheless the use of covariance matrices permits for more general transformations between Gaussian states which was employed in later studies, see e.g., [89, 90]. We will therefore stick with this approach in our presentation.
The quantum circuits generated by the unitary (31) can be rephrased in terms of the covariance matrix and displacement vector as follows
$\begin{split}{\bf G}=S{\bf G}S^{T},\quad{\bm{\omega}}=S{\bm{\omega}},\quad\text{where}\quad S\coloneqq e^{\mathbf{\Omega k}},\quad\end{split}$ (33)
Where ${\bf G},\bm{\omega}$ and ${\bf k}$ denote the covraiance matrix, displacement vector and symmetric matrix in the circuit (31) and the explicit indices have been suppressed to shorten the notation. The control functions $Y^{I}$ in equation (17) can be read as follows
@@
with $K_{I}$ is an orthonormal basis for the symplectic group. This then serves as the input for the complexity calculation with the cost assigned according to equation (25).
When evaluating complexity in QFTs, there are several key choices to be made. The initial studies [84, 85] focused on the simplest possible choices. The cost functions were selected according to equation (25), with $q_{I}=1$ and $p=1,2$. Only squeezing operations were considered in these early works. The target state was taken to be the regularized ground state of the bosonic field theory (30). Regularization was implemented via either a lattice discretization or a sharp momentum cutoff. The reference state was chosen as the ground state of the Hamiltonian
@@
where $\mu$ is a characteristic energy scale. This Hamiltonian describes a set of independent, or unentangled, oscillators with a characteristic frequency $\mu$. When constructing correlations at all scales, it is natural to begin with a state whose frequency is on the order of the inverse lattice spacing, so we define $\tilde{\mu}\coloneqq\mu\delta$, where $\delta$ is the lattice spacing.
@@
where for $p=1$ only an upper bound was found. In the previous equation $\mathcal{V}$ stands for the spatial volume of the system, $\delta$ is an ultraviolet cutoff (for example a lattice spacing) and the dots stand for less divergent terms. The full details of these calculations can be found in the original references and in the pedagogical review [2]. These results should be contrasted with those obtained from holography (114). We will later see that the $\mathcal{C}_{p=1}$ expression matches well with the volume-law divergence found in holographic calculations, see Eq. (114). Additionally, the choice of reference scale in the field theory parallels the choice of the counterterm scale $L_{ct}$ in holography, as we will see in, e.g., Equation (117a).
Subsequent studies explored a variety of alternative choices in the complexity calculations. These included different cost functions, free theories with other fields such as charged bosons, fermions, and gauge fields, weakly interacting theories and other target states, including the thermofield-double state, and mixed states [95, 89, 90, 88, 91, 92, 107, 93, 94, 108, 109, 110, 111, 112] . These studies concluded that, in general, the structure of divergences observed in QFT complexity calculations qualitatively matches the results from holography. This is not particularly surprising, since at very short distances (or equivalently, near the cutoff scale), the precise details of the theory tend to become irrelevant. On the other hand, the time dynamics of the theory shows significant deviations from holographic results, which is to be expected, as free and chaotic theories typically exhibit very different dynamical behavior. In particular, a natural expectation is that in free systems in finite volume upon penalization of Gaussian gates coupling distant parts complexity growth associated with Hamiltonian time evolution will terminate at times of the order of the volume (and even sooner without the aforementioned penalization [90]). In chaotic systems, one expects the complexity growth to occur over significantly longer timescales, exponential in the volume.
While many basic insights were obtained from the above approach, as mentioned earlier, it fails to reproduce many important properties of holographic complexity including those related to its temporal dynamics. This motivated a shift of focus to the the study of complexity in field theories more closely related to holography. This includes attempts to study complexity in conformal field theories focusing on universal conformal symmetry, which will be described in the following subsections.
@@ 5.5.2 Two dimensions - complexity of Virasoro circuits
Nielsen’s geometric approach is certainly very appealing from a mathematical (geometric) point of view, and it is tailor-made to mimic discrete quantum circuits built with gates available in our laboratories. However, defining gates in a generic QFT is not obvious and might generally involve considering all local operators, as an operator product can be represented as a sum over local operators within the operator product expansion. On top of that comes a question of choosing a meaningful cost function giving rise to an optimization problem that is not completely out of control.
These challenges forced the community to approach the problem of Nielsen’s complexity in QFT in a creative way. The progress made can be ultimately traced to originate from [30], which proposed to study Nielsen’s complexity of conformal transformations in 2-dimensional CFTs, i.e. coordinate transformations that preserve the form of the metric up to an overall scalar factor.
The immediate advantage of this approach was that it focused on universal properties shared by all CFTs, regardless of the value of the interaction strength. As a result, it brought a promise of finding precise gravitational realizations of complexity within holography. This goal was at least partially reached, as we will review in Sec. 7. Of course, this universality poses also a significant restriction, as resulting complexity measures will be insensitive to the microscopic dynamics with the results not allowing to distinguish between chaotic and integrable QFTs.
The connection with gates being local operators in a QFT occurs due to conformal transformations being generated by the most universal local operator present in any QFT, the energy momentum-tensor, smeared over a time-slice. For example, a familiar case of the conformal transformation is time translation generated by the Hamiltonian, which in turn is an integral over energy density - the time-time component of the energy-momentum tensor.
In more details, the setup of [30] and subsequent studies focused on in two-dimensional CFTs living on a Lorentzian cylinder. The spatial circle of unit radius will be parametrized by $\sigma$ and the physical time as $t$, see Fig. 6. In two dimensions the energy-momentum tensor of a CFT has two independent components that happen to commute with each other. In the following we will predominantly focus on one of them only, taken to be $T(\sigma)\equiv\frac{1}{2}T_{tt}+\frac{1}{2}T_{t\sigma}$. The algebraic structure underlying this construction is that of the Virasoro algebra
$[L_{n},L_{m}]=(n-m)L_{n+m}+\frac{c}{12}n(n^{2}-1)\delta_{n+m,0}\,.$ (37)
@@
In the above expressions $c$ is the central charge of the CFT, which can be thought of as measuring the number of degrees of freedom in the following sense: for $N$ free boson CFTs $c=N$, which is the number of bosonic degrees of freedom (whereas for $N$ free fermion CFTs, $c=\frac{N}{2}$). In holography, where the boundary CFTs are strongly coupled, $c\rightarrow\infty$. Since the algebra is the same for CFTs with the same value of $c$, one explicitly sees here the point we introduced earlier: the properties of Virasoro algebra-based complexity will not differ between $N\rightarrow\infty$ free bosons and strongly coupled holographic theories.
Figure 6: Two kinds of circuits one can consider in the context of conformal transformation in particular and QFT circuits more generally. In the latter situation one should think of $\sigma$ as a set of coordinates on spatial slices. Case a): The circuit parameter $\tau$ is an auxiliary parameter whose increase takes one CFT state to another CFT state. The transformation does not lead to a shift in the physical time $t$. This is the setting addressed in [30]. Case b): The circuit parameter $\tau$ is identified with the physical time $t$, as considered for the first time in [113]. Figure adapted from [113].
The setup of interest for [30] as well as subsequent works including [114, 115, 115, 116] are continuous unitary circuits
@@
In comparison with Eq. (17) discussed in the context of qubit systems, the integral over $\sigma$ is the analogue of the summation over the $I$-index, the $\epsilon(\tau)$ factor is the tangent space velocity $Y^{I}(t)$ and the generators can be thought of as $T(\sigma)$. Alternatively, one can use the Virasoro generators as defining circuit generators in which case summation over them is the analogue of $I$ in Eq. (17) and the tangent space velocity is related to $\epsilon_{n}(\tau)$.
In considering the complexity there is an important choice to be made regarding the interpretation of the $\tau$ parameter [113]. The first choice is to regard it as an auxiliary parameter with respect to the physical time $t$, see case (a) in Fig. 6. For this choice, relevant for [30], the function $\epsilon(\tau)$ associated with the diffeomorphism $\sigma\rightarrow f(\tau,\sigma)$ is given by
@@
The hardness minimization problem at hand concerns now finding a function of two parameters $f(\tau,\sigma)$ subject to the condition $f(\tau=0,\sigma)=\sigma$ and $f(\tau=1,\sigma)=f(\sigma)$, where $f(\sigma)$ is a desired diffeomorphisms that transforms the state $|h\rangle$ according to the CFT rules.
The papers [30, 114, 115, 117] uncovered and developed a very beautiful relation between optimizing cost functions
@@
and Euler-Arnold-type partial differential equations (PDEs) of significant relevance for mathematical physics. In particular, for the penalty schedules of the form
@@
with $a$ and $b$ being constants and $\delta(\sigma-\kappa)$ being the Dirac delta function, the solutions of the optimization problem are the solutions of the following paradigmatic PDEs [115]
  • 1.
@@
Furthermore, Refs. [115, 117] considered the Fubini-Study cost which corresponds to Eq. (43) with
@@
and solved the associated optimization in a perturbative manner for sample conformal transformation close to identity. This result was subsequently used for comparisons with holography, building on [118], see Sec. 7.
There are further generalizations of the approach described above. In [114] an additional global symmetry was incorporated connecting with the rich literature of the Kac-Moody groups. Furthermore, in [119] a scalar primary operator as an additional generator on top of the energy-momentum tensor operator was incorporated in a perturbative manner. The main limitation of the works prior to [119] is that the previously described unitary circuits remain bounded inside a conformal family (the so-called Verma module) of the Virasoro algebra, which is unrealistic. Results in [119] concern trajectories that move between different Verma modules in a given CFT and lessons on the dependence of the Fubini-Study cost of such circuits on the source fuction of the primary generator. This constitutes an important first step towards the general picture of Nielsen’s complexity with local operators viewed as generators advocated above.
Finally, another line of works devoted to complexity in CFT was concerned with circuits in which the physical time $t$ is the circuit parameter $\tau$ [113, 34, 116], see Fig. 6 case (b). In the context of the examples considered, this approach requires slightly different choice of the circuit velocity $\epsilon(t,\sigma)$. The main advantage of this approach is making the direct contact with the spacetime (path integral) formulation of QFTs. Since holography is naturally formulated in the language of path integrals, this led to the embedding of quantum circuits similar to the ones discussed in the present section on the boundary of the AdS spacetime and allow to find explicit gravitational duals to quantum circuits and the Fubini-Study cost function [34].

### v2

@@ 5.5 Relation between Nielsen complexity and entanglement: binding complexity
The penalty factors in the definition of Nielsen’s complexity can be attributed to the relative difficulty of applying certain generators. An interesting choice where analytic results can be obtained is the case in which our system is split into parts, and the penalty factors associated with generators acting within a given subsystem are much smaller than those of generators acting between multiple subsystems. This setting serves as a model to describe distributed quantum computing, where a large computer is divided into small quantum computers (called nodes). The operations between nodes, which are possibly located far away from each other, require significantly more resources than operations within each subsystem.
@@
The limiting case, where only operations acting between subsystems have non-vanishing cost, was referred to as binding complexity in [111]. There, the authors suggested that binding complexity quantifies the difficulty of distributing entanglement among multiple parties. The authors of Ref. [111] further suggested that the notion of binding complexity could serve as a measure for the robustness of entanglement. Here, one typically has in mind the $n$-party (with $n\geq 2$) generalizations of the W and GHZ states. The GHZ state is separable upon tracing out any subset of the parties, but the W state is not. In this sense, W states possess more robust entanglement. The authors of [111] argued that the binding complexity should scale as $\mathcal{O}(n)$ for GHZ states, and as $\mathcal{O}(n^{2})$ for $W$ states. The paper also presented results for the binding complexity of Gaussian states of free QFT. As we review next, we can get exact results for the binding complexity of spin chains in Nielsen’s framework when the cost of generators acting within each subsystem is small enough [115]. It turns out that cost functions satisfying this property relate Nielsen’s complexity to the Rényi min-entropy in a precise way. Moreover, the above cost functions are bounded by the entanglement entropy of a state over a subregion [121, 115].
@@
To illustrate these results, let us begin with a formal definition of the binding complexity. Consider splitting a quantum system (for instance, a spin chain) into $m$ subsystems $A_{k=1,2,\ldots,m}\,.$ The binding complexity is the minimal number of gates, each acting on a limited number of subsystems, needed to implement a target unitary with cost only incurred for gates acting between subsystems [111]. Below, we focus on 2-local gates which act on, at most, two subsystems each. Within Nielsen’s framework, binding complexity has a close analogue, obtained by choosing a particular cost function on the unitary space. To illustrate, consider a bipartite system whose Hamiltonian (17) decomposes as [^15] We denote the generators of the special unitary group in subsystem $A_{1}$ with index $a\in\{1,\dots,N_{A_{1}}^{2}-1\}$ from the beginning of the alphabet, and the generators in $A_{2}$ with index $i\in\{1,\dots,N_{A_{2}}^{2}-1\}$ from the middle part of the alphabet. We also denote with a sub(super)script $A_{1},A_{2}$ the subsystem where the local generators act.
@@
These penalties impose greater costs on entangling (or non-local) gates than on single-subsystem (or local) operations. Let us substitute the penalties (37) into the cost function (25) and perform the minimization (20) to obtain a complexity norm on the space of quantum states. Two simplifications follow: first, the space of states contains null directions, defined as the loci of states interconnected via local unitaries along which the cost function vanishes. Second, one can integrate out nondynamical degrees of freedom in the geodesic minimization. We obtain a norm $\mathcal{BF}[\lambda_{k},\dot{\lambda}_{k}]$ on the space of states. $\mathcal{BF}$ depends only on a state’s Schmidt coefficients $\lambda_{k}$ and their derivatives $\dot{\lambda}_{k}$.
We can analytically compute the binding state complexity for certain choices of the penalty factors $q_{ai}$ in Eq. (37), and of the parameter $p$ in Eq. (25). A Riemannian cost function is our first example. Let us write out the binding state complexity associated with the cost function (25), $p=2$, and the penalty factors (37) with $q_{ai}=1$ [115]:
@@
The relationship between binding complexity and Schmidt coefficients leads not only to exact results, but also bounds other quantum-information quantities. For instance, one can lower-bound the binding state complexity, in terms of the entanglement entropy of a state over a subregion $A_{1}$ [122, 123, 124, 121]. Denote by $S_{A_{1}}$ the entanglement entropy of subregion $A_{1}$, by $d$ the dimensionality of the smaller subsystem’s Hilbert space, and by $c$ a constant. The $p{=}1$ binding state complexity upper bounds $S_{A_{1}}$ as follows:
@@
Binding complexity also provides lower bounds for other notions of complexity, such as the geometrically local complexity [121]. Further details of the latter quantity will be discussed around Eq. (156).
Finally, let us mention that in the context of the holographic correspondence, binding complexity has been conjectured to be dual to the interior volume of multi-boundary wormholes [111].
@@ 5.6 Complexity in quantum field theories
As we will review in Sec. 7, Nielsen’s complexity experienced a revival following a conjectured relationship with gravitational quantities in the context of the holographic AdS/CFT correspondence. The holographic correspondence deals with Quantum Field Theories (QFTs) – many-body quantum systems which do not conserve particle number. To allow a quantitative comparison with holographic conjectures, it was therefore necessary to develop a method for calculating complexity in QFTs. Naturally, the first examples studied were those of free QFTs, first bosonic [99, 100] and then fermionic [103, 104], as reviewed in subsection 5.6.1. While these calculations are too simple for an exact comparison with chaotic theories in general – and with holographyic systems in particular – this approach helped to bridge the gap between heuristic arguments and more rigorous calculations. This framework also inspired later studies of complexity in more complicated field theories, including weakly interacting theories [106, 125], and conformal field theories, as described in subsections 5.6.2 and 5.6.3 below.
@@ 5.6.1 Free quantum field theories
The calculation of complexity in free field theories is based on Nielsen’s approach (see Sec. 5.1). In this framework, the evolution of quantum states is treated continuously, with a cost assigned to each possible evolution. To simplify the problem, one often restricts to unitary representations of finite-dimensional spaces. For free QFTs, this is typically achieved by restricting to transitions between Gaussian states. Such transitions are underpinned by the symplectic group for bosonic Gaussian states and the orthogonal group for fermionic Gaussian states.
Like many other quantities in QFT, complexity is expected to diverge because quantum correlations must be built at arbitrarily short distance scales. Therefore, a regulator should be introduced to handle short distances, or equivalently, large momenta. The exact result will depend on the choice of reference state, and in QFT, it is not immediately obvious which one to choose. Inspired by quantum computation, the reference state is often chosen to be completely unentangled, meaning that it lies outside the Hilbert space of the non-regulated theory, leading to the observed divergences.
@@
where $k_{(ab)}(t)$ is a symmetric matrix, move us between Gaussian states. A natural description of Gaussian states is in terms of their covariance matrix $G^{(ab)}$ and displacement vector $\omega^{a}$ defined using the first and second moments of the phase space variables
@@
where $\hat{\rho}$ is the Gaussian density matrix, and $\Omega^{[ab]}$ encodes the anti-symmetric commutation relations of the phase space variables. The covariance matrices were not the original formulation used to study complexity in bosonic QFT. Originally, the studies only focused on squeezing operations, and therefore expressing everything in terms of the wavefunction was sufficient. Nevertheless the use of covariance matrices permits for more general transformations between Gaussian states which was employed in later studies, see e.g., [104, 105]. We will therefore stick with this approach in our presentation.
The quantum circuits generated by the unitary transformation (41) can be recast in terms of their influence on the covariance matrix and displacement vector as follows
$\begin{split}{\bf G}=S{\bf G}S^{T},\quad{\bm{\omega}}=S{\bm{\omega}},\quad\text{where}\quad S\coloneqq e^{\mathbf{\Omega k}}\,.\quad\end{split}$ (43)
Here ${\bf G},\bm{\omega}$ and ${\bf k}$ denote the covariance matrix, displacement vector and symmetric matrix in the circuit (41) and explicit indices have been suppressed to shorten the notation. The control functions $Y^{I}$ in equation (17) can be read off as follows
@@
where $K_{I}$ is an orthonormal basis for the symplectic group. This setup is the input for the complexity calculation with the cost assigned according to equation (25).
When evaluating complexity in QFTs, there are several choices to be made. Initial studies [99, 100] selected cost functions according to equation (25), with $q_{I}=1$ and $p=1,2$, and only squeezing operations were considered. The target state was taken to be the regularized ground state of the bosonic field theory (40). Regularization was implemented either via a lattice discretization or a sharp momentum cutoff. The reference state was chosen as the ground state of the Hamiltonian
@@
where $\mu$ is a characteristic energy scale. This Hamiltonian describes a set of independent, or unentangled, oscillators with a characteristic frequency $\mu$. When constructing correlations at all scales, it is natural to begin with a state whose frequency is of the order of the inverse lattice spacing, so we define $\tilde{\mu}\coloneqq\mu\delta$, where $\delta$ is the lattice spacing.
@@
where for $p=1$ only an upper bound was found. In these equations $\mathcal{V}$ stands for the spatial volume of the system, $\delta$ is an ultraviolet cutoff (for example a lattice spacing) and the dots stand for less divergent terms (see the review [2]). These results should be contrasted with those obtained from holography (125). We will later see that the $\mathcal{C}_{p=1}$ expression matches well with the volume-law divergence found in holographic calculations, see Eq. (125). Additionally, the choice of reference scale in the field theory parallels the choice of the counterterm scale $L_{ct}$ in holography, as we will see in, e.g., Equation (128a).
Subsequent studies explored a variety of alternative cost functions, free theories with other fields such as charged bosons, fermions, and gauge fields, weakly interacting theories and other target states, including the thermofield-double state, and mixed states [111, 104, 105, 103, 106, 107, 125, 109, 110, 126, 127, 128, 129, 130]. These studies concluded that, in general, the structure of divergences observed in QFT complexity calculations qualitatively matches the results from holography. This is not particularly surprising, since at very short distances (or equivalently, near the cutoff scale), the precise details of the theory tend to become irrelevant. On the other hand, the time dynamics of the theory shows significant deviations from holographic results, which is to be expected, as free and chaotic theories typically exhibit very different dynamical behavior. In particular, a natural expectation is that in free systems of finite volume, after including penalties on Gaussian gates coupling distant parts of the system, complexity growth associated with Hamiltonian time evolution will terminate at times of the order of the volume (and even sooner without the penalties [105]). In chaotic systems, we expect complexity growth to occur over significantly longer timescales, exponential in the volume.
While many basic insights were obtained from the above approach, it fails to reproduce many important properties of holographic complexity, including those related to temporal dynamics. This observation motivated a shift of focus towards the study of complexity in field theories more closely related to holography. The latter included attempts to study complexity in conformal field theories via a focus on their universal conformal symmetry, which we describe below.
@@ 5.6.2 Two dimensions - complexity of Virasoro circuits
Nielsen’s geometric approach is appealing from a geometric point of view, and is tailor-made to mimic discrete quantum circuits built with gates available in laboratories. However, it is not clear how to define gates in a generic QFT. Indeed, we may need to consider all local operators, in view of the fact that an operator product can be represented as a sum over local operators via the operator product expansion. There is also the question of choosing a meaningful, yet tractable, cost function that gives rise to an optimization problem that is under control.
Such challenges forced the community to approach the problem of Nielsen’s complexity in QFT in a creative way. Progress can be traced back to [40], whose authors proposed to study Nielsen’s complexity for conformal transformations of 2-dimensional CFTs, i.e., coordinate transformations that preserve the form of the metric up to an overall (coordinate-dependent) scalar factor.
The immediate advantage of this approach is that it focuses on universal properties shared by all CFTs, regardless of the value of the interaction strength. As a result, it holds promise for understanding the gravitational dual of complexity within the AdS/CFT correspondence. Of course, this universality poses also a restriction, as the resulting complexity measures will be insensitive to the detailed microscopic dynamics.
The connection between gates and local operators in a QFT occurs because conformal transformations are generated by a local operator that is universally present in any QFT, the energy momentum-tensor, smeared over a time-slice. A familiar example of a conformal transformation is time translation generated by the Hamiltonian, which in turn is an integral of energy density - the time-time component of the energy-momentum tensor.
In more detail, [40] and subsequent studies focused on two-dimensional CFTs living on the Lorentzian cylinder. As illustrated in Fig. 6, we parametrize the spatial circle of unit radius by $\sigma$ and the physical time by $t$. In two dimensions, the energy-momentum tensor of a CFT has two independent components that commute with each other. In the following, we will predominantly focus on one of them only, $T(\sigma)\equiv\frac{1}{2}T_{tt}+\frac{1}{2}T_{t\sigma}$. The algebraic structure underlying this subsection’s construction is the Virasoro algebra
$[L_{n},L_{m}]=(n-m)L_{n+m}+\frac{c}{12}n(n^{2}-1)\delta_{n+m,0}\,,$ (47)
@@
In the above expressions, $c$ is the central charge of the CFT, which can be thought of as measuring the number of degrees of freedom in the following sense. For a CFT containing $N$ free bosons, $c=N$, which is the number of bosonic degrees of freedom; whereas for a CFT containing $N$ free fermions, $c=\frac{N}{2}$. In holography, where the boundary CFTs are strongly coupled, $c\rightarrow\infty$. Since the algebra is the same for all CFTs with the same value of $c$, one explicitly sees here the point we introduced earlier: the properties of Virasoro algebra-based complexity will not differ between $N\rightarrow\infty$ free bosons and strongly-coupled holographic theories.
Figure 6: Two kinds of circuits that can be considered in the context of conformal transformations, or, more generally, QFT (in this latter scenario, one should think of $\sigma$ as a set of coordinates on spatial slices). Case (a): The circuit parameter $\tau$ is an auxiliary parameter whose increase takes one CFT state to another CFT state. The transformation does not lead to a shift in the physical time $t$. This is the setting addressed in [40]. Case (b): The circuit parameter $\tau$ is identified with the physical time $t$, as considered for the first time in [131]. Figure adapted from [131].
The setup of interest for [40], as well as subsequent works including [132, 133, 133, 134], is given by continuous unitary circuits
@@
In comparison with Eq. (17) discussed in the context of qubit systems, the integral over $\sigma$ is the analogue of the summation over the $I$-index, the $\epsilon(\tau)$ factor is the tangent space velocity $Y^{I}(t)$, and the generators can be thought of as $T(\sigma)$. Alternatively, we can think of the Virasoro generators as defining circuit generators. In such case, the summation over them is the analogue of the sum over $I$ in Eq. (17), and the tangent space velocity is related to $\epsilon_{n}(\tau)$.
In considering the Nielsen’s complexity, there is an important choice to be made regarding the interpretation of the $\tau$ parameter [131]. The first choice is to regard it as an auxiliary parameter in relation to the physical time $t$, see case (a) in Fig. 6. For this choice, relevant for [40], the function $\epsilon(\tau)$ associated with the diffeomorphism $\sigma\rightarrow f(\tau,\sigma)$ is given by
@@
The complexity minimization problem at hand now requires us to find a function of two parameters $f(\tau,\sigma)$ subject to the condition $f(\tau=0,\sigma)=\sigma$ and $f(\tau=1,\sigma)=f(\sigma)$, where $f(\sigma)$ is a diffeomorphism that transforms the state $|h\rangle$ according to the CFT rules.
The papers [40, 132, 133, 135] uncovered and developed a very beautiful relation between optimizing cost functions of the form
@@
and Euler-Arnold-type partial differential equations (PDEs) of relevance for mathematical physics. In particular, for penalty schedules of the form
@@
with $a$ and $b$ constant and $\delta(\sigma-\kappa)$ the Dirac delta function, the solutions of the optimization problem are the solutions of the following paradigmatic PDEs [133]
@@
Furthermore, Refs. [133, 135] considered the Fubini-Study cost function, which corresponds to Eq. (53) with
@@
and solved the associated optimization perturbatively for sample conformal transformations close to identity. This result was subsequently used for comparisons with holography, building on [136], see Sec. 7.
There are further generalizations of the approach described above. Reference [132] connected with the rich literature of the Kac-Moody groups by incorporating an additional global symmetry. Furthermore, in Ref. [137] a scalar primary operator was added as an additional generator on top of the energy-momentum tensor operator. The main limitation of the works prior to [137] is that the previously described unitary circuits remain bounded inside a conformal family (the so-called Verma module) of the Virasoro algebra. Results in [137] concern trajectories that move between different Verma modules in a CFT, revealing the dependence of the Fubini-Study cost of such circuits on the source function of the primary generator. This investigation constitutes an important step towards arriving at a general picture of Nielsen’s complexity with local operators viewed as generators, as advocated above.
Finally, another line of works devoted to the study of complexity in CFT by examining trajectories in which the physical time $t$ is the circuit parameter $\tau$ [131, 44, 134], see Fig. 6 case (b). In the context of the examples considered, this approach requires slightly different choices of the circuit velocity $\epsilon(t,\sigma)$. The main advantage of this approach is that it makes direct contact with the spacetime (path integral) formulation of QFTs and with holography [44].